OpenStudy (anonymous):

if I have f(x)=x^3-ax+8 and they give f^-1(23)=3 how do i solve for a?

OpenStudy (amistre64):

or rather, find out when x=3 and y=23, what does a equal?

OpenStudy (amistre64):

23 =(3)^3-3a+8 15 =27 - 3a 5 =9 - a a = 4

OpenStudy (amistre64):

i get the right thought eventually ;)

OpenStudy (anonymous):

Omg it seems so simple now lol

OpenStudy (amistre64):

yep; just remember that an inverse is a relationship between x and y such that: f(x) = y ; f-1(y)=x

OpenStudy (anonymous):

thank you so much (: